### v2

# Robotic Policy Learning via Human-assisted Action Preference Optimization

Abstract
Establishing a reliable and iteratively refined robotic system is essential for deploying real-world applications. While Vision-Language-Action (VLA) models are widely recognized as the foundation model for such robotic deployment, their dependence on expert demonstrations hinders the crucial capabilities of correction and learning from failures. To mitigate this limitation, we introduce a Human-assisted Action Preference Optimization method named HAPO, designed to correct deployment failures and foster effective adaptation through preference alignment for VLA models. This method begins with a human-robot collaboration framework for reliable failure correction and interaction trajectory collection through human intervention. These human-intervention trajectories are further employed within the action preference optimization process, facilitating VLA models to mitigate failure action occurrences while enhancing corrective action adaptation. Specifically, we propose an adaptive reweighting algorithm to address the issues of irreversible interactions and token probability mismatch when introducing preference optimization into VLA models, facilitating model learning from binary desirability signals derived from interactions. Through combining these modules, our human-assisted action preference optimization method ensures reliable deployment and effective learning from failure for VLA models. The experiments conducted in simulation and real-world scenarios prove superior generalization and robustness of our framework across a variety of manipulation tasks.
[…]
1 Introduction
Ensuring safe interactions in unconstrained environments while fostering continuous improvement is crucial for the development of robust robotic manipulation systems in real-world scenarios [10, 26, 43]. Benefiting from the capacity for generalizable reasoning and scalable learning, Vision-Language-Action (VLA) models [3, 4, 8, 18, 41, 30] have been widely recognized as the foundation model for such robotic deployment systems. However, their performance in achieving high field-ready success rates in unconstrained, unpredictable real-world environments remains a significant limitation. This discrepancy presents a key challenge: how to integrate these developing Vision-Language-Action models into practical scenarios?
To ensure reliable deployment and continuous learning ability of robotic systems, interactive imitation learning frameworks [7, 17] have been developed to refine error-prone trajectories via iterative human-in-the-loop correction feedback. Among them, behavior cloning [15, 21, 33] has been widely utilized to fine-tune a base policy model with manually corrected intervention data in a supervised learning fashion. In contrast, recent methods [20, 24] seek to propose effective off-policy reinforcement learning algorithms from sub-optimal human intervention trajectories. However, behavior cloning fails to fully exploit failure trajectories, which are valuable signals for learning robust policies. At the same time, reinforcement learning methods encounter significant scalability limitations in training large-scale VLA models, due to the inherent instability and challenge of developing generalizable value functions. To date, the effective adaptation of VLA models for downstream manipulation tasks remains understudied, particularly within sub-optimal human intervention paradigms.
To enable reliable deployment and stable learning from interactions, we propose a Human-assisted Action Preference Optimization method named HAPO, for autoregressive VLA models. This method integrates two critical components: the human-robot collaboration framework for reliable deployment, and the action preference optimization process for iterative improvement of VLA models. As shown in Figure 1(a), the human-robot collaboration deployment framework allows real-time human interventions during policy execution, ensuring reliable task completion when the robot encounters challenging situations. To mitigate the proportion imbalance of corrective action, we propose a balanced sampling method to provide proportional representation of interaction data for further VLA preference optimization. As shown in Figure 1(b), we introduce the action preference optimization process to fully leverage the sub-optimal interaction trajectories for stable VLA models optimization, which helps avoid failure actions and encourages the adoption of corrective actions. Through iterative human-robot collaboration deployment and action preference optimization process, our method can continuously enhance the VLA model’s capabilities via environment interaction, ensuring sustained improvements in performance and adaptability to dynamic downstream manipulation tasks.
Figure 1: Pipeline of human-assisted action preference optimization method. This method consists of two components: (a) a human-robot collaboration deployment framework for reliable deployment and interaction trajectory collection with human intervention. (b) an action preference optimization process with adaptive reweighting for VLA models learning from sub-optimal human intervention trajectories. The size of circles represents their weight during training.
However, there exist two challenges in adopting preference alignment optimization on autoregressive VLA models fine-tuning: (1) Irreversible interaction: While LLMs require paired preference data for optimization, the irreversible interaction process makes it difficult to gather meaningful paired positive-negative action samples under the same observation conditions. (2) Token probability mismatch: Autoregressive VLA models discretize continuous actions into tokens, causing a mismatch between token probabilities and true prediction loss, complicating preference alignment in robot action tokens prediction. To address these problems, we first employ Kahneman & Tversky’s prospect theory [11, 38] to formulate a preference alignment objective function that learns from binary desirability signals derived from interaction. This objective relaxes the demands of preference pairs, making it particularly suitable for learning from irreversible robotic interaction trajectories. Furthermore, we propose an adaptive reweighting method that leverages decoded continuous actions to guide preference optimization in the discrete action token space. This approach addresses the challenge of action token probability mismatch via the dynamic modulation of sample-wise training weighting, thereby concentrating gradient optimization on failure-prone interaction actions. Through weight refinement, we apply preference alignment optimization to VLA models, enhancing performance and stability when learning from sub-optimal manipulation correction trajectories.
To systematically evaluate the effectiveness of our proposed system, we conduct comprehensive experiments in RoboMimic [27] simulation environments. The empirical results demonstrate that our HAPO method facilitates rapid adaptation in in-distribution scenarios while maintaining robust performance across a variety of unseen perturbations. Furthermore, lifelong learning experiments demonstrate the framework’s capacity for iterative improvement through human intervention. To evaluate the practical viability of the proposed framework, we conducted real-world experiments on fine-grained packing-hole tasks under a range of disruption conditions, demonstrating its robustness and applicability in real-world robotic manipulation scenarios.
2 Related Work
2.1 Vision-Language-Action Models
Achieving generalizable robotic manipulation remains a significant challenge within the field of robotics. Motivated by recent advances in foundation models [19, 37, 40, 42, 45], some works [6, 12, 39] attempt to construct large-scale real-world robotic datasets to facilitate the development of generalizable Vision-Language Action (VLA) models. Building upon these datasets, recent research [5, 18, 31] formulates robotic action prediction as a next token prediction problem within the framework of VLMs. In contrast, alternative studies [23, 36] investigate the applicability of diffusion-based methods to model multi-modal action distributions, thereby facilitating robustness in manipulation tasks. While these works focus on behavior cloning from expert demonstrations, Grape [46] proposes a trajectory-level preference alignment method to boost generalizability by incorporating both successful and failed trials. However, the requirements of paired trajectories under the same conditions make it infeasible in real-world scenarios. In this work, we build a human-assisted action preference optimization method to ensure reliable deployment in unconstrained robotic manipulation scenarios, alongside the preference alignment method that continuously refines the VLA model performance by systematically integrating human-in-the-loop intervention data.
2.2 Preference Alignment of Large Language Models
Contemporary methods [9, 28, 29, 47] implement Reinforcement Learning from Human Feedback (RLHF) through a two-stage method, which first trains a reward estimation model and optimizes LLMs to maximize the given estimated reward with a reinforcement learning method [13, 34]. However, this paradigm is slow and unstable in practice. DPO [32] proposes a single-step alternative that reparameterizes the RLHF objective into a closed-form loss function to directly maximize the log-likelihood margin between preferred and dispreferred outputs. Extending this framework, KTO [11] introduces the human-aware losses for learning from a binary signal of whether an output is desirable, which bypasses the need for intricate preference annotation altogether. In this work, we adapt the preference alignment optimization method for Vision-Language-Action models. Through an adaptive reweighting approach, we mitigate the irreversible interactions and token probability mismatch challenges when transferring preference learning methods from LLMs to VLA models.
2.3 Human-robot Interactive Learning
Interactive imitation learning [1, 7] has been proposed to refine robot actions through human feedback. While prior research [17, 21, 33] necessitates constant human supervision to intervene in the robot’s actions, more recent studies [14, 22, 44] have introduced dynamic models for automatic failure detection and real-time monitoring. In contrast, RLIF [24] leverages human intervention signals as rewards for off-policy RL, while HIL-SERL [25] presents a human-in-the-loop, vision-based RL system tailored for dexterous manipulation tasks. However, these RL approaches encounter difficulties in large-scale VLA model training, primarily due to unstable gradient optimization. In this work, we propose a human-assisted action preference optimization method to ensure the reliable deployment of robotics systems and the stable optimization of policies.
3 Human-assisted Action Preference Optimization
In this section, we introduce the human-assisted action preference optimization method as shown in algorithm 1 to enable reliable deployment and enhance continuous iterative improvement for the Vision-Language-Action (VLA) model. This framework comprises two components: a human-robot collaboration deployment framework to support real-time human intervention for reliable task execution, and an action preference alignment optimization process with adaptive reweighting for stable autoregressive VLA models fine-tuning through human interventions.
3.1 Human-robot Collaboration Deployment
We first collect an expert demonstration dataset $\mathcal{D}_{e}=\{\tau_{e}^{i}\}_{i=1}^{i=N}$, where each trajectory $\tau_{e}^{i}$ consists of observation-action pairs with expert annotations: $\tau_{e}^{i}=\{(o_{t}^{i},a_{t}^{i},c_{t}^{i})\}_{t=1}^{t=T}$, where $c_{t}^{i}=1$ indicates that $a_{t}^{i}$ is executed by human expert. We employ behavior cloning to fine-tune the pretrained VLA model on these expert demonstrations, obtaining an initial base policy $\pi_{\theta}^{0}$. This policy is then deployed for interaction trajectory collection.
During policy execution, the human operator monitors policy execution and intervenes when the policy encounters challenging scenarios. Through this process, we could collect a set of interaction trajectories $\mathcal{D}_{h}=\{\tau^{i}_{h}\}_{i=i}^{i=M}$, where $c_{t}^{i}=2$ represents the action is corrected by human intervention while $c^{i}_{t}=1$ denotes the action is executed by policy. Further, we re-label the interaction trajectories to categorize the actions taken in the $K$ steps preceding human interventions as undesirable, annotated with $c_{t}^{i}=0$. For each trajectory, we discretize the continuous action $a$ into discrete action token $\hat{a}$. Finally, we combine the expert demonstrations $\mathcal{D}_{e}$ and the interaction dataset $\mathcal{D}_{h}$ for further robotic action preference optimization.
3.2 Action Preference Optimization
To maximize the utility of sub-optimal interaction trajectories and ensure stable fine-tuning of the VLA model, we adopt the preference alignment optimization method to guide the model to learn from corrections and avoid failures.
Although previous Reinforcement Learning with Human Feedback (RLHF) methods [2, 32] have proven effective in LLM fine-tuning, there are additional challenges for the VLA models preference optimization in robotic manipulation:
  • The irreversible robotic manipulation process makes it challenging to acquire meaningful paired positive-negative actions under the same observational conditions.
  • The mapping of continuous robotic actions to discrete tokens by autoregressive VLAs causes a mismatch between token probability and continuous action errors, complicating preference optimization in action token prediction.
Algorithm 1 Human-assisted Action Preference Optimization
1:Notations:
2:$\mathcal{D}_{e}$: expert demonstrations, $\mathcal{D}_{h}$: interaction dataset, $\pi$: interaction policy
3:Warm-start phase
4:Collect $\mathcal{D}_{e}\leftarrow\{\tau_{1}^{e},\dots,\tau_{N}^{e}\}$
5:Initialize BC policy $\pi^{0}_{\theta}$
6:$\theta^{*}\leftarrow\arg\max_{\theta}\mathbb{E}_{(o,a)\sim\mathcal{D}^{e}}\left[\log\pi^{0}_{\theta}(a|o)\right]$
7:$\mathcal{D}_{h}^{0}\xleftarrow{}\mathcal{D}_{e}$
8:Deployment-optimization loop
9:for $i\leftarrow 0$ to $X$ do $\pi_{ref}\leftarrow\pi_{\theta}^{i}$ $D_{h}^{i+1}\leftarrow\textsc{Deployment}(\pi^{i}_{\theta},\mathcal{D}_{h}^{i})$ $\pi^{i+1}_{\theta}\leftarrow\textsc{Optimization}(\pi_{\theta}^{i},\pi_{ref},\mathcal{D}_{h}^{i+1})$
10:end for
11:
12:function Deployment($\pi_{\theta},\mathcal{D}_{h}$)
13:     for n interaction rollouts do
14:         while task does not succeed do
15:              if human intervenes then
16:                  $a_{t}\xleftarrow{}human,c_{t}\xleftarrow{}2$
17:                  $c_{t-K:t-1}\leftarrow 0$
18:              else
19:                  $a_{t}\xleftarrow{}\pi_{\theta}(o_{t}),c_{t}\xleftarrow{}1$
20:              end if
21:              $\tau_{i}^{h}\xleftarrow{}\tau^{h}_{i}\cup(o_{t},a_{t},c_{t})$
22:         end while
23:         $\mathcal{D}_{h}\xleftarrow{}\mathcal{D}_{h}\cup\tau_{i}^{h}$
24:     end for
25:     return $\mathcal{D}_{h}$
26:end function
27:function Optimization($\pi_{\theta}$,$\pi_{ref}$,$\mathcal{D}_{h}$) 28:     for n gradient steps do 29:         Balanced Sample $(o_{i},a_{i},c_{i})_{i=1}^{i=B}\sim\mathcal{D}_{h}$ 30:         $l_{i}\xleftarrow{}|\pi_{\theta}(o_{i})-a_{i}|_{1}$ 31:         $w_{i}\xleftarrow{}\frac{l_{i}}{\sum_{i=1}^{i=B}l_{i}}$ 32:         if $c_{i}\neq 0$ then $\lambda_{D_{i}}=1-e^{-\beta_{D}*w_{i}}$ 33:         else $\lambda_{U_{i}}=e^{-\beta_{U}*w_{i}}$ 34:         end if 35:         $\theta^{*}=\arg\max_{\theta}\mathbb{E}[-v(o,a,\pi_{\theta},\pi_{ref},\lambda_{D},\lambda_{U})]$ 36:     end for 37:     return $\pi_{\theta^{*}}$ 38:end function
To address these issues, we adopt Kahneman & Tversky’s prospect theory [38] for preference alignment optimization with binary desirability signals and propose an adaptive reweighting method to bridge the gap between discrete token prediction and continuous action regression. We first estimate the reward function $r_{\theta}$ of our model $\pi_{\theta}$ as standard approach [29, 32, 35]:
$r_{\theta}(o,\hat{a})=\log\frac{\pi_{\theta}(\hat{a}|o)}{\pi_{\text{ref}}(\hat{a}|o)},$ (1)
where $\hat{a}$ is the discrete action token and the reference model $\pi_{ref}$ is the base model $\pi_{\theta}^{i}$ at the beginning of each deployment-optimization loop shown in algorithm 1. Following [11, 38], we formulate the utility function $v$ as below to estimate the relative gain on the robotic data:
$v(o,\hat{a})=\begin{cases}\lambda_{D}\sigma\left(r_{\theta}(o,\hat{a})-z_{0}\right)&\text{if }\hat{a}\sim\hat{a}_{\text{desirable}}\\ \lambda_{U}\sigma\left(z_{0}-r_{\theta}(o,\hat{a})\right)&\text{if }\hat{a}\sim\hat{a}_{\text{undesirable}},\end{cases}$ (2)
where $\lambda_{D}$ and $\lambda_{U}$ are utilized for importance sampling, the $\sigma$ is the sigmoid function. To ensure that the model $\pi_{\theta}$ does not deviate excessively from the reference model $\pi_{ref}$, a penalty term $z_{0}$ is introduced. This term is defined as the KL-Divergence between $\pi_{\theta}$ and $\pi_{ref}$: $z_{0}=KL(\pi_{\theta}||\pi_{ref})$. Incorporating $z_{0}$ into the optimization process guides the model to learn from preference pair data while simultaneously preserving knowledge acquired from prior models. We employ the following loss function $L$ to optimize the model $\pi_{\theta}$ using preference optimization with desirability signals:
$L(\pi_{\theta},\pi_{ref})=\mathbb{E}_{x,y\sim D^{h}}[-v(x,y)].$ (3)
By minimizing the loss function, we aim for the model $\pi_{\theta}$ to get higher rewards for desirable pairs while avoiding predicting undesirable actions, in comparison to the reference model $\pi_{ref}$.
However, directly applying the preference alignment optimization from LLMs to autoregressive VLA models is problematic, primarily due to the differences in their respective token definitions. While word tokens correspond to distinct subwords, action tokens necessitate a non-differentiable mapping to continuous ground truth actions. This creates a discrepancy between the token classification probabilities and the regression loss associated with the continuous robotic actions.
To bridge the gap between token classification and continuous action regression in autoregressive VLA models, we introduce an adaptive reweighting method. This approach guides the model to prioritize samples exhibiting large regression errors by first estimating the L1 loss of the continuous action $l$ for each sample, followed by batch-level normalization as detailed below:
$w_{i}=\frac{l_{i}}{\sum_{i=1}^{i=B}l_{i}}.$ (4)
The normalized weighting scheme operates by: 1) for desirable data, increasing the weight of samples with high action prediction errors, and 2) for undesirable data, increasing the weight of samples whose actions are proximate to the failure actions. By adaptively adjusting the values of $\lambda_{D}$ and $\lambda_{U}$ in Equation 7 using the normalized weights, we gain fine-grained control over the relative influence of each sample during training:
$\lambda_{D}$ $=1-e^{-\beta_{D}*w},$ (5)
$\lambda_{U}$ $=e^{-\beta_{U}*w}.$ (6)
By incorporating preference alignment optimization via sample-wise weight refinement, we enhance the performance and optimization stability of the VLA model, when learning from sub-optimal manipulation correction trajectories.
In conclusion, we propose a human-assisted action preference optimization method as demonstrated in algorithm 1. This approach leverages the human-robot collaboration deployment for reliable task execution and interaction trajectories collection, while the action preference optimization process provides stable autoregressive VLA optimization with adaptive reweighting. Through iterative human-robot collaboration deployment and action preference optimization, we could achieve continual improvement from interaction with environments for autoregressive VLA models.
4 Experiments
To comprehensively evaluate our human-assisted action preference optimization method for effective downstream adaptation, we propose experiments to validate the following questions:
  • Could our method promote effective adaptation to in-distribution scenarios? Section 4.2
  • Could our method perform effectively in novel scenarios with various disruptions? Section 4.3
  • Could our method achieve iterative improvement during deployment? Section 4.4
  • Could our method generalize to various VLA models? Scetion 4.5
  • Could our method be applied in fine-grained real-world scenarios? Section 4.6
4.1 Experiment Settings
Implementation Details. In this work, we fine-tune the OpenVLA [39] model for target manipulation tasks as the base model. We employ LoRA [16] for parameter-efficient tuning, configuring rank $r=32$ with a batch size of 16 across 8 NVIDIA A100 GPUs. Further, we deploy the base model to interact with environments, where human operators perform real-time corrective interventions via a SpaceMouse device to rectify failures during execution. Based on the interaction trajectories collected during task execution, we fine-tune the base model $\pi_{ref}$ with our action preference optimization method, using a learning rate of $-55e{\text{-}}$ 5 and a batch size of 8 across 4 NVIDIA A100 GPUs. We set $K=10$ to automatically identify and annotate undesirable behaviors. To ensure the stability of preference alignment training, we employ balanced sampling to ensure that each batch contains 50% expert actions, 25% human intervention actions, and 25% failure actions.
Simulation Environments Details. We conduct experiments on RoboMimic [27] simulation. In the RoboMimic environment, we fine-tune the pretrained OpenVLA model for 4 long-horizon manipulation tasks with 300 expert demonstrations. To optimize policy with human preferences, we collect 50 trajectories per task under different seeds for RoboMimic tasks. For evaluation, we conduct 50 trials under three unseen seeds for each task, and report the average success rate.
Table 1: Comparison experiment results across 4 manipulation tasks in RoboMimic Simulation. The results demonstrate that our adaptive reweighting preference optimization method achieves stable improvement compared with other behavior cloning and preference optimization methods.
Methods Coffee_D0 StackThree_D0 ThreePieceAssebly_D0 Square_D0 Mean
Base policy 44% 46% 44% 28% 40.5%
Dagger [33] 42% 50% 36% 28% 39.0%
Sirius [21] 34% 52% 34% 38% 39.5%
DPO [32] 52% 46% 28% 22% 37.0%
TPO [46] 54% 54% 40% 18% 41.5%
KTO[11] 48% 52% 46% 32% 43.5%
Ours 60% 54% 46% 32% 48%
4.2 Comparison Experiments
We compare our adaptive reweighting preference optimization method with other approaches to evaluate the effectiveness for VLA model fine-tuning. To ensure fairness, we fine-tune OpenVLA [18] for manipulation tasks as a base model and improve the base model with other comparison methods.
  • Dagger [33]: We mix the expert demonstrations with interaction trajectories, fine-tuning the base model using a behavior cloning objective.
  • Sirius [21]: we apply sample reweighting to prioritize human intervention data and fine-tune the base model using a weighted behavior cloning loss.
  • DPO [32]: We generate paired negative samples for interaction trajectories by perturbing the actions predicted from the base model with Gaussian noise, and fine-tune the base model using these paired data with the DPO method.
  • TPO [46]: We select positive and negative samples in interaction trajectories based on the timing of intervention, then fine-tune model with trajectory-wise preference optimization.
  • KTO [11]: We select positive and negative samples in interaction trajectories based on the timing of intervention. Then, we sample positive and negative trajectories and optimize the base model with KTO, with the constraint $z_{0}=KL(\pi_{\theta}||\pi_{ref})$.
As shown in Table 1, we first compare the behavior cloning objective methods. The results reveal that after fine-tuning with interaction data, these methods fail to outperform the base model, which demonstrates that existing behavior cloning approaches struggle to achieve efficient adaptation in large-scale VLA models. A key challenge stems from the distribution shift between expert trajectories and interaction trajectories. Without mechanisms to retain the base model’s knowledge under the standard behavior cloning objectives, this shift makes it particularly difficult for large-scale VLA models to fit the multimodal distribution of expert and interaction datasets.
We further provide results of the preference optimization based methods. By integrating a regularization constraint with the reference model $\pi_{ref}$, these methods could maintain useful knowledge from the reference model while achieving improvement from interaction trajectories.
Among all compared preference learning based methods, DPO yields the weakest performance. This result stems from its exclusive reliance on synthetic paired failure actions for optimization, which lacks exposure to real-world errors essential for teaching robots mistake avoidance through interaction. On the other side, TPO fails to deliver stable performance gains on multiple tasks while our attains stable performance gains relative to the base model. The TPO method employs negative samples to regularize model preference alignment optimization, but introduces instability through random sampling. In contrast, our method utilizes KL divergence to estimate the mean margin between the updated model and the reference model, which not only enables more stable learning but also better preserves prior knowledge. Compared with KTO, our method leverages the adaptive reweighting method to achieve more precise control over the importance weights of both positive and negative samples, delivering more notable performance improvements.
Figure 2: In the position disruption setting, we change the position of the stick from a fixed point to a random position from the rectangle in the Square_D0 task as illustrated in (a). In the background disruption setting, we replace the background with the gray one in the StackThree_D0 task as shown in (b). In the texture disruption setting, we replace the red blocks with the wooden ones.
Table 2: The results on disruption scenarios.
Methods Pos Dis. Bg Dis. Tex Dis. Mean
Base policy 12% 42% 10% 21.3%
Dagger 18% 46% 4% 22.7%
Sirius 12% 42% 2% 18.7%
DPO 14% 26% 2% 14.0%
TPO 18% 32% 8% 19.3%
KTO 20% 46% 6% 24.0%
Ours 26% 46% 12% 28.0%
Table 3: The results on original tasks.
Methods Square StackThree ThreePiece Mean
Base policy 28% 46% 44% 39.3%
Dagger 16% 46% 30% 30.7%
Sirius 18% 48% 18% 28%
DPO 20% 50% 30% 33.3%
TPO 30% 36% 40% 35.3%
KTO 30% 46% 42% 39.3%
Ours 34% 62% 40% 45.3%
4.3 Generalization to Novel Tasks
In this section, we assess our framework’s generalization capability under three novel scenarios, as illustrated in Figure 2. (1) Position Disruption: For the Square_D0 task, we replace the fixed initial stick position with randomized placements within a bounded operational area. (2) Background Disruption: In the StackThree_D0 task, we substitute the default white background with a gray one. (3) Texture Disruption: In the ThreePiece_D0 task, the original red blocks are transmuted to wood-grain visual properties. These experiments systematically evaluate robustness against spatial, background, and visual texture variations. To fine-tune the base model on novel disruption scenarios, we collect 20 interaction trajectories under disruption scenarios and combine them with 20 expert demonstrations from the original task for subsequent fine-tuning.
Our objective is to develop a human-assisted action preference optimization method that facilitates continuous improvement, enabling performance enhancements in novel disruption scenarios while retaining original task capabilities during model fine-tuning. Thus, we evaluate the performance of the fine-tuned model across both disruption scenarios and original scenarios.
As shown in Tabel 4.2, the base policy exhibits some degree of performance degradation in disruption scenarios. However, the performance decline is relatively minor in cases of background disruption, whereas disruptions in object texture and position significantly impact performance. Both behavior cloning methods and preference optimization methods struggle to achieve significant performance improvements in novel disruption scenarios. In contrast, our approach can effectively adapt to new disruption scenarios through adaptive reweighting.
Table  4.2 presents the performance of the optimized model after being fine-tuned from disruption data on the original task. The results reveal that behavior cloning methods exhibit severe catastrophic forgetting, resulting in substantial performance degradation. By contrast, the preference optimization method achieves mitigated performance decline with the constraints of the reference model.
Besides, our approach utilizes adaptive reweighting to effectively integrate knowledge from both expert demonstrations and interaction trajectories. This mechanism not only facilitates learning from diverse data sources but also leads to improved performance on the original task.
4.4 The Performance of Lifelong Learning
To investigate whether our method can iteratively improve via environment interaction, we deploy our model to interact with environments while updating the model every 20 interaction rollouts. Additionally, we provide comparison results using a behavior cloning policy trained with the same number of expert demonstrations as our baseline. For each updated model, we conduct 50 trials and report the success rate.
As shown in Figure 3(a-b), our method achieves superior performance compared to the baseline, demonstrating its ability to effectively leverage sub-optimal human intervention trajectories for iterative model improvement. When the base policy exhibits diminishing improvement with increasing expert demonstrations, our method enables continual performance gains from the interaction trajectories. Besides, this improvement trend is accompanied by a corresponding reduction in the required human intervention ratio, as shown in Figure 3(c)
Figure 3: Lifelong learning results of our mehtod.
4.5 Generalization to various VLA models
To validate that our method can be adapted to different VLA models, we applied our approach to fine-tune the $\pi 0$-FAST [31] model. $\pi 0$-FAST applies discrete cosine transform encoding to encode the action chunking into discrete tokens for VLA training. To adopt this model for downstream tasks, We regenerate the action tokenizer with 5 action chunking step for each task.
As shown in Table 9, the base model could achieve a higher success rate compared with the OpenVLA model, benefiting from its ability to predict action chunking for robotic manipulation. Further, we compare our method with both the behavior cloning method and preference optimization method, the results demonstrate that our method could achieve consistent improvement for the $\pi 0$-FAST fine-tuning. The results prove that our method could our method can be applied to the fine-tuning of various VLA models, achieving consistent performance gains.
Table 4: The results on $\pi 0$-FAST model.
Methods Coffee_D0 StackThree_D0
Base policy 68% 64%
Dagger 64% 66%
TPO 48% 52%
Ours 76% 74%
Table 5: The results on real-world experiments.
Methods In Dis. Pos Dis. Bg Dis. Tex Dis.
Base policy 65% 25% 10% 25%
Dagger 65% 10% 10% 25%
TPO 75% 40% 20% 45%
Ours 85% 55% 30% 55%
4.6 Real-world Experiments
In this work, we conduct the challenging fine-grained robotic manipulation task “Insert the square into the stick” as shown in Figure 7(a), which requires the robot to grasp the square and precisely insert into the stick. To collect expert demonstrations, we utilize the spacemouse device to gather 100 high-quality trajectories at an action frequency of 20 Hz. We fine-tune the OpenVLA model with the collected demonstration as the base model. Further, we deploy the base model to interact with environments and propose the real-time human-in-the-loop interventions to collect 20 interaction trajectories for subsequent preference alignment optimization. To fairly compare performance, all methods are evaluated under the same experimental setup, and we report the average success rate from 20 trials.
To comprehensively evaluate our method, experiments are conducted not only under in-distribution but also across three distinct disruption settings as shown in Figure 7(b-e): (1) Position Disruption: We change the position distribution of the stick. (2) Background Disruption: We replace the tablecloth from white to brown. (3) Texture Disruption: We replace the black stick to the green one.
As demonstrated in Table 4.5, our method demonstrated robust adaptability to these downstream disruption scenarios. The results empirically validate the method’s practical utility for real-world deployment in unstructured environments.
Figure 4: Demonstrations of real-world experiments with disruption settings.
5 Conclusion
In this work, we introduce the human-assisted action preference optimization method, consisting of two critical components: a human-robot collaboration framework for reliable deployment and an action preference optimization process with adaptive reweighting for stable VLA model optimization. Through our method, we could promote continuous improvement during the deployment of VLA models. We hope our method could bring insights for efficient and effective VLA model adaptation on downstream manipulation tasks.
Limitation. While our work study the preference alignment optimization for VLA models, the experiments are based solely on autoregressive VLA models. Future work should explore a broader range of VLA frameworks, including regression-based approaches and diffusion policy models, to ensure the generalizability of our method across different architectures.